OpenStudy (anonymous):

an airplane leaves an airport and flies due west 170 miles and then 180 miles in the direction 190˚10'. assuming the earth is flat, how far is the plane from the airport at this time(to the nearest mile)?

OpenStudy (therealmeeeee):

ok

OpenStudy (therealmeeeee):

Loading more…

OpenStudy (therealmeeeee):

isnt it c^2 = a^2 +b^2

OpenStudy (therealmeeeee):

here, the angle C is 39.17+90, and u have a, b, only c left to find, which is the answer

OpenStudy (therealmeeeee):

replace them in a2 + b2 - 2ab cos(C) = c2 to find c

OpenStudy (therealmeeeee):

120 mi150 mi39.17

OpenStudy (dumbcow):

 use Law of cosines to find "x"